# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v3

@@ Appendix F Bounding Total Variation Distance
As $|\Delta|\to n$, we prove that the total variation (TV) distance between the empirical distribution of $z_{\Delta}$ and $z_{[n]}$ goes to 00. Here, recall that $z_{i}\in\mathcal{Z}=\mathcal{S}_{l}\times\mathcal{A}_{l}$, and $z_{\Delta}=\{z_{i}:i\in\Delta\}$ for $\Delta\in{[n]\choose k}$. Before bounding the total variation distance between $F_{z_{\Delta}}$ and $F_{z_{\Delta^{\prime}}}$, we first introduce Lemma C.5 of Anand and Qu [2024] which can be viewed as a generalization of the Dvoretzky-Kiefer-Wolfowitz concentration inequality, for sampling without replacement. We first make an important remark.
@@
First, observe that if $\Delta$ is an independent random variable uniformly supported on (${[n]\choose k}$, then $s_{\Delta}$ and $a_{\Delta}$ are also independent random variables uniformly supported on the global state (${s_{[n]}\choose k}$ and the global action (${a_{[n]}\choose k}$. To see this, let $\psi_{1}:[n]\to\mathcal{S}_{l}$ where $\psi_{1}(i)=s_{i}$ and $\xi_{1}:[n]\to\mathcal{A}_{l}$ where $\xi_{1}(i)=a_{i}$. This naturally extends to $\psi_{k}:[n]^{k}\to\mathcal{S}_{l}^{k}$, where $\psi_{k}(i_{1},\dots,i_{k})=(s_{i_{1}},\dots,s_{i_{k}})$ and $\xi_{k}:[n]^{k}\to\mathcal{A}_{l}^{k}$, where $\xi_{k}(i_{1},\dots,i_{k})=(a_{i_{1}},\dots,a_{i_{k}})$ for all $k\in[n]$. Then, the independence of $\Delta$ implies the independence of the generated $\sigma$-algebra. Further, $\psi_{k}$ and $\xi_{k}$ (which are a Lebesgue measurable function of a $\sigma$-algebra) are sub-algebras, implying that $s_{\Delta}$ and $a_{\Delta}$ must also be independent random variables.
@@
Recall that $z:=(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}$. From Lemma F.3, substituting $\mathcal{B}_{l}=\mathcal{Z}_{l}$ yields:
@@
We now present an alternate bound for the total variation distance, where the distance actually goes to 00 as $|\Delta|\to n$. For this, we use the fact that the total variation distance between two product distributions is subadditive.
@@
Thus, from Equation 27:
@@
From Lemma F.6, setting $\Delta=[n]$ also recovers $\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})=0$.
@@
From combining the total variation distance bound in Lemma F.4 and the Lipschitz continuity bound in Theorem E.3 with $\sum_{t=0}^{T}\gamma^{T}\leq\frac{1}{1-\gamma}$ for $\gamma\in(0,1)$, we have:
@@ Appendix G Using the Performance Difference Lemma
where ${\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{[n]})=\hat{\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{u},F_{s_{\Delta\setminus u}})$ is the global agent’s action and ${\pi}_{k,m}^{i,\mathrm{est}}(s_{g},s_{[n]}):=\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Here, $\Delta_{i}$ is a random variable supported on (${[n]\setminus i\choose k-1}$, $u$ is a random variable uniformly distributed on $[n]$, and $\Delta$ is a random variable uniformly distributed on $[n]\setminus u$. Then, denote the optimal policy $\pi^{*}$ given by
@@
Lemma G.4 shows a uniform bound on
@@
(independent of $\Delta^{i},\forall i\in[n]$), allowing the sums and the counts in the denominator will cancel out. Observe that $\hat{\pi}_{k,m}^{*}:\mathcal{S}\to\mathcal{A}$ and $\pi^{*}:\mathcal{S}\to\mathcal{A}$ are deterministic functions. Therefore, denote $a=\pi^{*}(s)$. Then, from Lemma G.6,
@@
Similarly, from Corollary G.7, we have that
@@
which yields the claim. We defer parameter optimization to Lemma G.8.∎
@@
Fix $s\in\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$. For each $j\in[n]$, suppose we are given $T$-length sequences of random variables $\{\Delta^{j}_{1},\dots,\Delta^{j}_{T}\}_{j\in[n]}$, distributed uniformly over the support (${[n]\setminus j\choose k-1}$. Further, suppose we are given a fixed sequence $\delta_{1},\dots,\delta_{T}$, where each $\delta_{t}\in(0,1]$ for $t\in[T]$. Let
@@
where the $\epsilon_{k,m}$ follows from Lemma 4.3. From Theorem F.8, we have that with probability at least $1-\delta_{t}$,
@@
For the first term in the first inequality, we use $\mathbb{E}[X\mathbbm{1}\{X\leq c\}]\leq c$. For the second term, we trivially bound $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\hat{Q}_{k}^{*}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})$ by the maximum value $Q^{*}$ can take, which is $\frac{\tilde{r}}{1-\gamma}$ by Lemma C.4. In the second inequality, we use the fact that the expectation of an indicator function is the conditional probability of the underlying event. The second inequality follows from Lemma G.5.
@@
Changing the notation of the policy to ${\pi}_{k,m}^{\mathrm{est}}$ in the proofs of Lemmas G.6 and G.5, such that $a\sim\tilde{\pi}_{k,m}^{\mathrm{est}}$ then yields that:
@@
Setting $\delta=\frac{(1-\gamma)^{2}}{2\tilde{r}\varepsilon|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}$ in the bound for the optimality gap in Theorem G.3 gives:
@@
Finally, using the probabilistic bound of $\epsilon_{k,m}\leq\tilde{O}(\frac{1}{\sqrt{k}})$ from Lemma G.10 yields that with probability at least $1-\frac{1}{100e^{k}}$,
@@ G.1 Bounding the Bellman Error
This section is devoted to the proof of Lemma 4.3.
@@
If $m\in\mathbb{N}$ is the number of samples in the Bellman update, there exists a universal constant $20<c_{0}<$ 2 and a Bellman noise $0<\epsilon_{k,m}\leq\frac{1}{1-\gamma}$ such that
@@
where $\mathcal{B}_{t}$ denotes the event that all $(s,a)\in\mathcal{S}\times\mathcal{A}$ have been visited at least once between time 00 and time $t$, and $\mathbb{P}(\mathcal{B}_{t}|s_{0},a_{0})$ denotes the probability of $\mathcal{B}_{t}$ conditioned on the initial state $(s_{0},a_{0})$.
@@
If $T=\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$, SUB-SAMPLE-MFQ: Learning runs in time $\tilde{O}(T|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{3.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\tilde{r})$, while accruing a Bellman noise $\epsilon_{k,m}\leq\tilde{O}(1/\sqrt{k})$ with probability at least $1-\frac{1}{100e^{k}}$, .
@@
Since $T=\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}>\frac{1}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$, the condition holds and $\|\hat{Q}_{k}^{T}-\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{1}{\sqrt{k}}$. Then, rearranging Equation 30 and incorporating the convergence error of the $\hat{Q}_{k}$-function, one has that with probability at least $1-\rho$,
@@
Substituting this in Equation 32 yields that with probability $1-\rho$,
@@
$O(mT|\mathcal{S}_{g}||\mathcal{S}_{l}||\mathcal{A}_{g}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|})=\tilde{O}(|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{3.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\tilde{r}),$
@@ Appendix H Generalization to Stochastic Rewards
0:  A multi-agent system as described in Section 2. Parameter $T$ for the number of iterations in the initial value iteration step. Sampling parameters $k\in[n]$ and $m\in\mathbb{N}$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$.
@@
When ${\pi}^{\mathrm{est}}_{k,m}$ is derived from the randomized empirical value iteration operation and applying our online subsampling execution in Algorithm 2, we get
@@
Then, applying the triangle inequality to $\epsilon_{k,m}$ allows us to derive a probabilistic bound on the optimality gap between $\hat{Q}_{k,m}^{\mathrm{est}}$ and $Q^{*}$, where the gap is increased by $\frac{1}{\sqrt{200k}}$, and where the randomness is over the stochasticity of the rewards. Then, the optimality gap between $V^{\pi*}$ and $V^{{\pi}_{k,m}^{\mathrm{est}}}$, for when the policy $\hat{\pi}^{\mathrm{est}}_{k,m}$ is learned using Algorithm 4 in the presence of stochastic rewards obeying H.1 follows
@@
First, note that that through the naïve averaging argument, the optimality gap above still decays to 00 with probability decaying to $1$, as $k$ increases.
@@ Appendix I Partially Relaxing the Offline Learning assumption: Off-Policy Learning
A limitation of the planning Algorithm 1 is that it learns $\hat{Q}_{k}^{*}$ in an offline manner by assuming a generative oracle access to the transition functions $\mathbb{P}_{g},\mathbb{P}_{l}$, and reward function $r(\cdot,\cdot)$. In certain realistic RL applications, such a generative oracle might not exist, and it is more desirable to perform off-policy learning where the agent continues to learn in an offline manner but from historical data Fujimoto et al. [2019]. In this setting, the agents learn the target policy $\hat{\pi}_{k}^{*}$ using data generated by a different behavior policy (the strategy it uses to explore the environment). There is a significant body of work on the theoretical guarantees in off-policy learning Chen et al. [2021b], Chen and Maguluri [2022b], Chen et al. [2021a, 2025]. In fact, these previous results are amenable to transforming guarantees about offline $Q$-learning to off-policy $Q$-learning, at the cost of $\log|\mathcal{S}_{g}||\mathcal{S}_{l}|^{k}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{k}$ factors in the runtime. Therefore, this section is devoted to showing that our previous result satisfy the assumptions of transforming offline $Q$-learning to off-policy $Q$-learning for the subsampled $\hat{Q}_{k}$-function, and we further show that, in expectation, can maintain the decaying optimality gap of $\tilde{O}(1/\sqrt{k})$ of the learned policy $\pi_{k}$, where the randomness is over the heuristic exploration policy $\pi_{b}$.
@@
Note that the update in Equation 42 does not include an expectation and can be computed in a single trajectory via historical data. We make the following assumption:
@@
Let $\pi_{k}$ be the policy learned through off-policy $\hat{Q}_{k}$-learning. Then, under I.1, we have that that with probability at least $1-\frac{1}{100e^{k}}$,
@@
    $\|\hat{Q}_{k}(F_{S_{\Delta}},F_{A_{\Delta}})-\hat{Q}_{k}(F_{S^{\prime}_{\Delta}},F_{A_{\Delta}})\|\leq\frac{1}{1-\gamma}\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\|F_{S^{\prime}_{\Delta}}-F_{S_{\Delta}}\|_{1}$, for any $S_{\Delta},S^{\prime}_{\Delta}\in\mathcal{S}_{g}\times\mathcal{A}_{l}^{k}$ and $A_{\Delta}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ by Theorem E.3,
@@
    $\|\hat{Q}_{k}\|\leq\frac{\tilde{r}}{1-\gamma}$ by Lemma C.4,
@@
    $\|\hat{Q}^{t+1}_{k}(S_{\Delta},A_{\Delta})-\tilde{Q}_{k}^{t+1}(S_{\Delta},A_{\Delta})\|_{\infty}\leq\gamma\|\hat{Q}_{k}^{t}\ -\tilde{Q}_{k}^{t}(S_{\Delta},A_{\Delta})\|_{\infty}$ by Lemma C.6,
@@
    The Markov chain $\mathcal{M}_{S}$ enjoys a rapid mixing property from I.1,
@@
Suppose $\alpha_{t}=\alpha$ for all $t\geq 0$, where $\alpha$ is chosen such that $\alpha t_{\alpha}(\mathcal{M}_{S_{\Delta}})\leq c_{Q,0}\frac{(1-\beta_{1})^{2}}{\log|\mathcal{S}_{g}||\mathcal{S}_{l}^{k}||\mathcal{A}_{g}|\mathcal{A}_{l}^{k}|}$, where $c_{Q,0}$ is a numerical constant. Then, under Proposition I.3, for all $t\geq t_{\alpha}(\mathcal{M}_{S_{\Delta}})$, we have
@@
With this sample complexity, by the triangle inequality we also recover an expected-value analog of Theorem F.8.
@@
In turn, following the argument in the proof of Theorem G.3, it is straightforward to verify that this leads to a result on the expected performance difference using off-policy learning:
@@
Suppose $\mathcal{S}_{g}\subset\mathbb{R}^{\sigma_{g}},\mathcal{S}_{l}\subset\mathbb{R}^{\sigma_{l}},\mathcal{A}_{g}\subset\mathbb{R}^{\alpha_{g}},\mathcal{A}_{l}\subset\mathbb{R}^{\alpha_{l}}$ are bounded compact sets. From the Baire category theorem, the underlying field $\mathbb{R}$ can be replaced to any set of Baire first category, which satisfies the property that there exists a dense open subset. In particular, replace 2.1 with a boundedness assumption.
@@
:-$\coloneq\langle\bar{\phi}_{n}(s,a),\bar{\mu}_{n}(s^{\prime})\rangle:$

### v4

@@ Appendix F Bounding Total Variation Distance
As $|\Delta|\to n$, we prove that the total variation (TV) distance between the empirical distribution of $z_{\Delta}$ and $z_{[n]}$ goes to 0. Here, recall that $z_{i}\in\mathcal{Z}=\mathcal{S}_{l}\times\mathcal{A}_{l}$, and $z_{\Delta}=\{z_{i}:i\in\Delta\}$ for $\Delta\in{[n]\choose k}$. Before bounding the total variation distance between $F_{z_{\Delta}}$ and $F_{z_{\Delta^{\prime}}}$, we first introduce Lemma C.5 of Anand and Qu [2024] which can be viewed as a generalization of the Dvoretzky-Kiefer-Wolfowitz concentration inequality, for sampling without replacement. We first make an important remark.
@@
First, observe that if $\Delta$ is an independent random variable uniformly supported on $n]\choose k$}, then $s_{\Delta}$ and $a_{\Delta}$ are also independent random variables uniformly supported on the global state $s_{[n]}\choose k$} and the global action $a_{[n]}\choose k$}. To see this, let $\psi_{1}:[n]\to\mathcal{S}_{l}$ where $\psi_{1}(i)=s_{i}$ and $\xi_{1}:[n]\to\mathcal{A}_{l}$ where $\xi_{1}(i)=a_{i}$. This naturally extends to $\psi_{k}:[n]^{k}\to\mathcal{S}_{l}^{k}$, where $\psi_{k}(i_{1},\dots,i_{k})=(s_{i_{1}},\dots,s_{i_{k}})$ and $\xi_{k}:[n]^{k}\to\mathcal{A}_{l}^{k}$, where $\xi_{k}(i_{1},\dots,i_{k})=(a_{i_{1}},\dots,a_{i_{k}})$ for all $k\in[n]$. Then, the independence of $\Delta$ implies the independence of the generated $\sigma$-algebra. Further, $\psi_{k}$ and $\xi_{k}$ (which are a Lebesgue measurable function of a $\sigma$-algebra) are sub-algebras, implying that $s_{\Delta}$ and $a_{\Delta}$ must also be independent random variables.
@@
Recall that $z:=(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}$. From Lemma˜F.3, substituting $\mathcal{B}_{l}=\mathcal{Z}_{l}$ yields:
@@
We now present an alternate bound for the total variation distance, where the distance actually goes to 0 as $|\Delta|\to n$. For this, we use the fact that the total variation distance between two product distributions is subadditive.
@@
Thus, from Equation˜29:
@@
From Lemma˜F.6, setting $\Delta=[n]$ also recovers $\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})=0$.
@@
From combining the total variation distance bound in Lemma˜F.4 and the Lipschitz continuity bound in Theorem˜E.3 with $\sum_{t=0}^{T}\gamma^{T}\leq\frac{1}{1-\gamma}$ for $\gamma\in(0,1)$, we have:
@@ Appendix G Using the Performance Difference Lemma
where ${\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{[n]})=\hat{\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{u},F_{s_{\Delta\setminus u}})$ is the global agent’s action and ${\pi}_{k,m}^{i,\mathrm{est}}(s_{g},s_{[n]}):=\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Here, $\Delta_{i}$ is a random variable supported on $n]\setminus i\choose k-1$}, $u$ is a random variable uniformly distributed on $[n]$, and $\Delta$ is a random variable uniformly distributed on $n]\setminus u$. Then, denote the optimal policy $\pi^{*}$ given by
@@
Lemma˜G.4 shows a uniform bound on
@@
(independent of $\Delta^{i},\forall i\in[n]$), allowing the sums and the counts in the denominator will cancel out. Observe that $\hat{\pi}_{k,m}^{*}:\mathcal{S}\to\mathcal{A}$ and $\pi^{*}:\mathcal{S}\to\mathcal{A}$ are deterministic functions. Therefore, denote $a=\pi^{*}(s)$. Then, from Lemma˜G.6,
@@
Similarly, from Corollary˜G.7, we have that
@@
which yields the claim. We defer parameter optimization to Lemma˜G.8.∎
@@
Fix $s\in\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$. For each $j\in[n]$, suppose we are given $T$-length sequences of random variables $\{\Delta^{j}_{1},\dots,\Delta^{j}_{T}\}_{j\in[n]}$, distributed uniformly over the support $n]\setminus j\choose k-1$}. Further, suppose we are given a fixed sequence $\delta_{1},\dots,\delta_{T}$, where each $\delta_{t}\in(0,1]$ for $t\in[T]$. Let
@@
where the $\epsilon_{k,m}$ follows from Lemma˜4.3. From Theorem˜F.8, we have that with probability at least $1-\delta_{t}$,
@@
For the first term in the first inequality, we use $\mathbb{E}[X\mathbbm{1}\{X\leq c\}]\leq c$. For the second term, we trivially bound $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\hat{Q}_{k}^{*}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})$ by the maximum value $Q^{*}$ can take, which is $\frac{\tilde{r}}{1-\gamma}$ by Lemma˜C.4. In the second inequality, we use the fact that the expectation of an indicator function is the conditional probability of the underlying event. The second inequality follows from Lemma˜G.5.
@@
Changing the notation of the policy to ${\pi}_{k,m}^{\mathrm{est}}$ in the proofs of Lemmas˜G.6 and G.5, such that $a\sim\tilde{\pi}_{k,m}^{\mathrm{est}}$ then yields that:
@@
Setting $\delta=\frac{(1-\gamma)^{2}}{2\tilde{r}\varepsilon|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}$ in the bound for the optimality gap in Theorem˜G.3 gives:
@@
Finally, using the probabilistic bound of $\epsilon_{k,m}\leq\tilde{O}(\frac{1}{\sqrt{k}})$ from Lemma˜G.10 yields that with probability at least $1-\frac{1}{100e^{k}}$,
@@ G.1 Bounding the Bellman Error
This section is devoted to the proof of Lemma˜4.3.
@@
If $m\in\mathbb{N}$ is the number of samples in the Bellman update, there exists a universal constant $0<c_{0}<2$ and a Bellman noise $0<\epsilon_{k,m}\leq\frac{1}{1-\gamma}$ such that
@@
where $\mathcal{B}_{t}$ denotes the event that all $(s,a)\in\mathcal{S}\times\mathcal{A}$ have been visited at least once between time 0 and time $t$, and $\mathbb{P}(\mathcal{B}_{t}|s_{0},a_{0})$ denotes the probability of $\mathcal{B}_{t}$ conditioned on the initial state $(s_{0},a_{0})$.
@@
[…]
@@
Since $T=\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}>\frac{1}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$, the condition holds and $\|\hat{Q}_{k}^{T}-\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{1}{\sqrt{k}}$. Then, rearranging Equation˜32 and incorporating the convergence error of the $\hat{Q}_{k}$-function, one has that with probability at least $1-\rho$,
@@
Substituting this in Equation˜34 yields that with probability $1-\rho$,
@@
[…]
@@ Appendix H Generalization to Stochastic Rewards
0: A multi-agent system as described in Section˜2. Parameter $T$ for the number of iterations in the initial value iteration step. Sampling parameters $k\in[n]$ and $m\in\mathbb{N}$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$.
@@
When ${\pi}^{\mathrm{est}}_{k,m}$ is derived from the randomized empirical value iteration operation and applying our online subsampling execution in Algorithm˜2, we get
@@
Then, applying the triangle inequality to $\epsilon_{k,m}$ allows us to derive a probabilistic bound on the optimality gap between $\hat{Q}_{k,m}^{\mathrm{est}}$ and $Q^{*}$, where the gap is increased by $\frac{1}{\sqrt{200k}}$, and where the randomness is over the stochasticity of the rewards. Then, the optimality gap between $V^{\pi*}$ and $V^{{\pi}_{k,m}^{\mathrm{est}}}$, for when the policy $\hat{\pi}^{\mathrm{est}}_{k,m}$ is learned using Algorithm˜4 in the presence of stochastic rewards obeying ˜H.1 follows
@@
First, note that that through the naïve averaging argument, the optimality gap above still decays to 0 with probability decaying to $1$, as $k$ increases.
@@ Appendix I Partially Relaxing the Offline Learning assumption: Off-Policy Learning
A limitation of the planning Algorithm˜1 is that it learns $\hat{Q}_{k}^{*}$ in an offline manner by assuming a generative oracle access to the transition functions $\mathbb{P}_{g},\mathbb{P}_{l}$, and reward function $r(\cdot,\cdot)$. In certain realistic RL applications, such a generative oracle might not exist, and it is more desirable to perform off-policy learning where the agent continues to learn in an offline manner but from historical data Fujimoto et al. [2019]. In this setting, the agents learn the target policy $\hat{\pi}_{k}^{*}$ using data generated by a different behavior policy (the strategy it uses to explore the environment). There is a significant body of work on the theoretical guarantees in off-policy learning Chen et al. [2021b], Chen and Maguluri [2022b], Chen et al. [2021a, 2025].
In fact, these previous results are amenable to transforming guarantees about offline $Q$-learning to off-policy $Q$-learning, at the cost of $\log|\mathcal{S}_{g}||\mathcal{S}_{l}|^{k}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{k}$ factors in the runtime. Therefore, this section is devoted to showing that our previous result satisfy the assumptions of transforming offline $Q$-learning to off-policy $Q$-learning for the subsampled $\hat{Q}_{k}$-function, and we further show that, in expectation, can maintain the decaying optimality gap of $\tilde{O}(1/\sqrt{k})$ of the learned policy $\pi_{k}$, where the randomness is over the heuristic exploration policy $\pi_{b}$.
@@
Note that the update in Equation˜44 does not include an expectation and can be computed in a single trajectory via historical data. We make the following ergodicity assumption:
@@
Let $\pi_{k}$ be the policy learned through off-policy $\hat{Q}_{k}$-learning. Then, under ˜I.1, we have that that with probability at least $1-\frac{1}{100e^{k}}$,
@@
    $\|\hat{Q}_{k}(F_{S_{\Delta}},F_{A_{\Delta}})-\hat{Q}_{k}(F_{S^{\prime}_{\Delta}},F_{A_{\Delta}})\|\leq\frac{1}{1-\gamma}\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\|F_{S^{\prime}_{\Delta}}-F_{S_{\Delta}}\|_{1}$, for any $S_{\Delta},S^{\prime}_{\Delta}\in\mathcal{S}_{g}\times\mathcal{A}_{l}^{k}$ and $A_{\Delta}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ by Theorem˜E.3,
@@
    $\|\hat{Q}_{k}\|\leq\frac{\tilde{r}}{1-\gamma}$ by Lemma˜C.4,
@@
    $\|\hat{Q}^{t+1}_{k}(S_{\Delta},A_{\Delta})-\tilde{Q}_{k}^{t+1}(S_{\Delta},A_{\Delta})\|_{\infty}\leq\gamma\|\hat{Q}_{k}^{t}\ -\tilde{Q}_{k}^{t}(S_{\Delta},A_{\Delta})\|_{\infty}$ by Lemma˜C.6,
@@
    The Markov chain $\mathcal{M}_{S}$ enjoys a rapid mixing property from ˜I.1,
@@
Suppose $\alpha_{t}=\alpha$ for all $t\geq 0$, where $\alpha$ is chosen such that $\alpha t_{\alpha}(\mathcal{M}_{S_{\Delta}})\leq c_{Q,0}\frac{(1-\beta_{1})^{2}}{\log|\mathcal{S}_{g}||\mathcal{S}_{l}^{k}||\mathcal{A}_{g}|\mathcal{A}_{l}^{k}|}$, where $c_{Q,0}$ is a numerical constant. Then, under Proposition˜I.3, for all $t\geq t_{\alpha}(\mathcal{M}_{S_{\Delta}})$, we have
@@
With this sample complexity, by the triangle inequality we also recover an expected-value analog of Theorem˜F.8.
@@
In turn, following the argument in the proof of Theorem˜G.3, it is straightforward to verify that this leads to a result on the expected performance difference using off-policy learning:
@@
Suppose $\mathcal{S}_{g}\subset\mathbb{R}^{\sigma_{g}},\mathcal{S}_{l}\subset\mathbb{R}^{\sigma_{l}},\mathcal{A}_{g}\subset\mathbb{R}^{\alpha_{g}},\mathcal{A}_{l}\subset\mathbb{R}^{\alpha_{l}}$ are bounded compact sets. From the Baire category theorem, the underlying field $\mathbb{R}$ can be replaced to any set of Baire first category, which satisfies the property that there exists a dense open subset. In particular, replace ˜2.1 with a boundedness assumption.
@@
Example J.7 (Federated Learning).
Consider a peer-to-peer learning setting [Chaudhari et al., 2024] with $n$ agents, where each agent possesses a common neural network architecture $f(\mathbf{x};\mathbf{\theta}):\mathbb{R}^{N}\to\mathbb{R}^{M}$, parameterized by $\theta\in\mathbb{R}^{P}$. Here, each local agent has a common local objective function $L_{q}(\mathbf{\theta}):\mathbb{R}^{P}\to\mathbb{R}$ defined using a local dataset $\mathcal{D}_{q}\stackrel{{\scriptstyle\Delta}}{{=}}\{x_{q_{i}},y_{q,i}\}_{i=1}^{D}$. For instance, $L_{q}(\mathbf{\theta})$ can be the local empirical risk function that evaluates the performance of $f(\mathbf{x};\mathbf{\theta})$ on a task involving the local dataset $\mathcal{D}_{q}$, where the agents cooperate to learn the optimal $\mathbf{\theta}^{*}$ that minimizes the global objective, which is a weighted average of all local objectives. Formally, the agents collaboratively aim to find $\mathbf{\theta}^{*}$ satisfying:
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{P}}L(\theta)$
$L(\theta)=\frac{1}{Q}\sum_{q=1}^{Q}L_{q}(\theta)\stackrel{{\scriptstyle\Delta}}{{=}}\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{D}\sum_{i=1}^{D}\ell(f(x_{q,i};\theta),y_{q,i})$
Here, the global agent may act to balance the loss among agents, by ensuring that the variance on the losses across the agents is small. If the variance on the losses of the agents is large, it could assign a large cost to the system, and signal that the policies of the local agents must favor the mean state, promoting convergence.
@@
$\coloneq\langle\bar{\phi}_{n}(s,a),\bar{\mu}_{n}(s^{\prime})\rangle$